Consider the following information on returns and probabilities:

  State Probability    X   Z       

Boom .25 15% 10%

Normal .60 10% 9%

Recession .15 5% 10%

What is the standard deviation for a portfolio with an investment of $6,000 in asset X and $4,000 in asset Z?

A) 6.10%

B) 8.15%

C) 1.85%

D) 3.00%

E) 5.18%
Answer #1

Investment in asset X = $ 6000

Investment in asset Z = $ 4000

Total investment = $ 6000 + $ 4000 = $ 10000

Probability in boom (PB) = 0.25

Probability in Normal (PN) = 0.60

Probability in Recession (PR) = 0.15

Solution :-

weight of asset X (Wx) 6000 – 0.6 - O. 6 looco Weight of Wr) asset ? Yooo - = 0.4 10000 Portfolio deturn in Boom (PIRB) - Ret- Portfolio detuan in Recession (PLR) p) Return of x in recession x W x + [ Return of 2 in recession & W₂] = (5%)(0.6) + (10%60 Sot = 10.029430.25 +1 -0.0046) 0.60 + -0.0306) *0.15 SO = (0,00086436) 0.25 +(6.0002)16) 0.60 +(0.00093636) 0.15 SD = 0.00